Tammy wants to complete a jigsaw puzzle. On the first day, she manages to fit 20% of the puzzle. On the second day, she fits another 60 pieces of the puzzle. As a result, the ratio of the number of fitted pieces to the number of unfitted pieces becomes 2:5. How many pieces does the jigsaw puzzle consist?
Rmks: Pls try not to use algebra.

Answer by nerdybill(7384)

Let x = number of pieces in the puzzle
.20x + 60 = pieces fitted on second day
.
From: "the ratio of the number of fitted pieces to the number of unfitted pieces becomes 2:5" we have:
(.20x + 60)/(x-(.20x + 60)) = 2/5
(.20x + 60)/(x-.20x - 60) = 2/5
(.20x + 60)/(.80x - 60) = 2/5
5(.20x + 60) = 2(.80x - 60)
x + 300 = 1.6x - 120
420 = .6x
700 = x